OpenStudy (anonymous):

Nakim simplified 3 times the square root of 2x plus x times the square root of 8x minus 5 times the square root of 18x and got -10x times the square root of 2x for an answer. Part 1: Using complete sentences, explain what Nakim did wrong. Part 2: Show all your work to simplify the expression. (You can use "sqrt()" to show a square root. For example, 7 times the square root of a can be written as 7sqrt(a). )

OpenStudy (anonymous):

help please :(

OpenStudy (anonymous):

\[3\sqrt{2x} + x \sqrt{8x} -5\sqrt{18x} \neq -10x \sqrt{2x}\] Note that one term has a xsqrt(8x) while the other terms don't have it. that term was originally \[\sqrt{8x ^{2}x}\]

OpenStudy (anonymous):

so what did he do wrong?

OpenStudy (anonymous):

As for the correct solution.. I will get to the last step.\[3\sqrt{2x} + x \sqrt{4 \times 2x} - 5\sqrt{9 \times 2x}\] than you get \[3\sqrt{2x} + 4x \sqrt{2x} - 5(3)\sqrt{2x}\]

OpenStudy (anonymous):

I pointed out the mistake you have to explain it.

OpenStudy (anonymous):

Is x^2 the same as x ?

OpenStudy (anonymous):

2x?

OpenStudy (anonymous):

excuse me.. Did you try to do the full solution? Take it from where I stopped.

OpenStudy (anonymous):

alright hold on I'm doing it

OpenStudy (anonymous):

I have no clue :(

OpenStudy (anonymous):

alright pay close attention pretend \[\sqrt{2x} = y\] so now we get this \[3y + 4xy - 15y\] can you simplify this.

OpenStudy (anonymous):

common dude lets go step it up!!

OpenStudy (anonymous):

-12y+4xy

OpenStudy (anonymous):

Alright can you factor a y?

OpenStudy (anonymous):

can you factor a 4?

OpenStudy (anonymous):

how do i do that?

OpenStudy (anonymous):

just divide both expressions by 4y

OpenStudy (anonymous):

-3y+x?

OpenStudy (anonymous):

\[4y(x-3)\]

OpenStudy (anonymous):

so is that the answer?

OpenStudy (anonymous):

Yea dude just substitute the sqrt(2x) back for y

OpenStudy (anonymous):

thanks dude

OpenStudy (anonymous):

yw =))